Session: Feb 1998 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Sergio Mariotti at Tue Feb 17 08:44:07 MET 1998) Minor failures
The experiment has been recorded in Thick tapes with the Control Head Table prepared for Thin Tapes.

Torun (K.Borkowski at Tue Feb 17 09:31:27 MET 1998) Success
No known problems

Jodrell (Alastair Gunn at Tue Feb 17 11:25:10 MET 1998) Minor failures
Telescope parked due to high winds until about 17:05 UT. No other known problems.

Cambridge (Alastair Gunn at Tue Feb 17 11:25:48 MET 1998) Minor failures
Vacuum lost at 02:46, repositioned at 03:00 but lost again, repositioned at 03:08. Probably lost a pass here. Vacuum set at 15 inches because of some tape flying at double speed on our VLBA recorder. No other known problems.

Onsala (A. Polatidis at Tue Feb 17 16:47:10 MET 1998) Success
Observed with the old LCP polariser.

Noto (C.Trigilio at Fri Feb 20 15:40:33 MET 1998) Success

Effelsberg (R. Schulze at Tue Feb 24 08:38:59 MET 1998) Minor failures
We had problems in the beginning, lost keyboard input before the experiment start, restarted, after schedule call found that program had locked onto scan 2. So we moved tape fast forward and started with pass 2. We do not understand why that happened. According to the operator recording at the end they lost everything after 07:09:40

Medicina ( at Thu Feb 26 13:08:50 MET 1998) Minor failures

Calibration (jean\-francois Desmurs at Mon Mar 2 18:46:06 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb98 directory under the name em028.antab. Read the comments in the file to get further information.